Given the following table, use the matrix method to derive the constant and slope parameters of the sample regression function: Marks = f(Weekly party hours). X (Weekly party hours) 2,4,6,8,10,12,14,16 Y (Marks) 80,77,71,67,66,58,49,36


1
Expert's answer
2021-11-08T17:25:40-0500

We first tabulate the values of X and Y and calculate "X^2" and "XY"



Using matrix;matrix


"X'X=\begin{bmatrix}
 8 & 72 \\
 72 & 816
\end{bmatrix}"


"X'Y=" "\begin{bmatrix}
 504 \\
 4048
\end{bmatrix}"


"X'X^{-1}=\begin{bmatrix}
 0.607 & -0.0536\\
 -0.0536 & 0.00595
\end{bmatrix}"


b="(X'X^{-1})X'Y"


="\begin{bmatrix}
 0.607 & -0.0536\\
 -0.0536 & 0.00595
\end{bmatrix}""\begin{bmatrix}
 504 \\
 4048
\end{bmatrix}"


="\begin{bmatrix}
 88.9552\\
-2.9288
\end{bmatrix}"


constant=88.9552

slope=-2.9288